# The Psychology of Learning from Machines: Anthropomorphic AI and the Paradox of Automation in Education

### v1

@@ III-C Transferability and Limitations
Although our dataset derives from philosophical rather than technical content, underlying cognitive mechanisms—trust calibration, anthropomorphic projection, and temporal evolution—are content-independent. Engineering students engaging with AI-generated tutorials, design critiques, or code explanations must similarly decide when to trust AI outputs and how to balance efficiency against deep understanding.
@@ IV-A4 Connection to Dual Bias Literature
Although our data come from general AI-generated debates rather than tutoring contexts, the patterns mirror well-documented dual biases in human–automation interaction. Learners simultaneously exhibit automation bias (overtrust) and algorithm aversion (rejecting systems after errors)[dietvorst2015algorithm]. Students accept algorithmic guidance for objective tasks like math or coding but grow skeptical when judgments are subjective or value-laden[castelo2019task], a pattern that intensifies as stakes rise [horowitz2024bending]. The expertise paradox further complicates reliance: novices overuse AI due to low confidence, while experts distrust it when confident in their own judgment [gaube2020experienced]. EPFL students even rated identical feedback lower once told it came from AI [kaser2024ai]. Our findings align with these dynamics: audiences embraced AI for abstract philosophical content (analogous to structured technical material) but reacted with strong polarization to moral-ethical topics (analogous to design reasoning or safety-critical decisions). This supports differentiated deployment in engineering education—AI tutors for well-defined theoretical tasks, human facilitation for open-ended, ethical, or judgment-intensive domains where algorithm aversion may be protective. The central challenge is cultivating appropriate reliance, avoiding both uncritical acceptance and premature rejection, through trust-calibration instruction, metacognitive training on when to verify outputs, and assessments that measure independent reasoning rather than AI-assisted performance.
@@ IV-B1 Behavioral Profiling Reveals Distinct Engagement Styles
Our behavioral profiling of commenters identified 73% as “Emotional Broadcasters”—users writing extensive commentary (100+ characters) with high affective language. This contrasts sharply with engineering tutorial audiences, where 70% were classified as “Casual Observers” posting brief, low-affect comments.
@@
Here, personal connection directly impacts academic performance—the content serves immediate practical needs.
@@ IV-B2 Parasocial Dynamics Distinguish Human from AI
This comment treats AI as an analytical tool rather than a teaching partner. The absence of gratitude, relationship-building language, and future engagement requests suggests audiences implicitly recognize AI as fundamentally different from human educators—different in kind, not merely in implementation.
@@
This ironic contrast humanizes some AIs while characterizing others as amusingly mercenary—treating AIs as moral agents with distinct personalities, a phenomenon not observed in engineering contexts.
@@ IV-B3 Connection to Anthropomorphism Research
These observed patterns align with established research showing that social responses to technology are automatic reactions to cues persisting despite intellectual understanding of machine nature [reeves1996media]—a phenomenon our data suggests occurs even in asynchronous contexts without direct interaction. Epley’s three-factor theory [epley2007seeing] may explain the intensity variations we observed: media hype about AI “intelligence” (elicited agent knowledge) primes audiences to overestimate capabilities, while the contrasting engagement styles—extensive philosophical commentary (89.4 words) for AI debates versus brief technical questions (9.2 words) for engineering tutorials—suggest different motivational drivers (sociality motivation versus effectance motivation for problem-solving).
However, research documents concerning manifestations: a study of 56 students using Betty’s Brain with physical robot Cozmo found that while physical presence increased initial enjoyment, higher perceived sociability negatively predicted task performance ($\beta=-0.28$, $p<0.05$) through distraction effects, with students of lower language proficiency finding AI tutors more disturbing [bettysbrain]—suggesting anthropomorphic design must carefully account for cognitive resource constraints.
Our findings suggest an anthropomorphism design challenge: features that may enhance initial engagement (conversational warmth, personality traits) could distract from learning or create inappropriate attachments. The absence of parasocial behaviors toward AI in our data—despite extensive engagement—may indicate that audiences implicitly recognize appropriate boundaries when consuming AI-generated content asynchronously. However, this may not hold for interactive AI tutoring contexts where students seek help during moments of frustration or confusion.
For engineering education, this suggests calibrated anthropomorphism based on task demands. When teaching complex technical material requiring sustained focus, minimizing anthropomorphic cues that consume cognitive resources may be warranted. For scaffolding motivation during problem-solving, some warmth may prove beneficial. Critical is awareness that vulnerable students—experiencing loneliness, struggling academically, or from underrepresented backgrounds—may be particularly susceptible to inappropriate attachment.
@@ IV-C2 Trust Evaluation Showed Implicit Acceptance
Explicit trust evaluation was rare in both domains, with 96.7% (n=145) of AI comments and 90.7% (n=136) of engineering comments exhibiting “neutral stance”—neither questioning nor affirming credibility ($\chi^{2}=10.23$, $p=0.037$, Cramér’s $V=0.185$). This suggests implicit trust in educational YouTube content regardless of creator type.
@@ IV-C3 Connection to Learning and Expertise Development
Our temporal patterns resonate with long-standing concerns about how deep learning develops. Weinberg argued that genuine understanding requires extended temporal exposure rather than instantaneous mastery [weinberg1971psychology], and our finding that analytical engagement emerges primarily after a week while superficial reactions dominate early periods suggests a similar dynamic. If deep engagement unfolds only with time, AI tutors optimized for instant answers risk shortchanging the sustained practice through which expert intuition develops.
This aligns with Bainbridge’s automation ironies [bainbridge1983ironies]: systems designed to support cognition may unintentionally erode skills through disuse. Recent work demonstrates these patterns persist with generative AI [simkute2025genai], with mathematical AI dependency negatively correlating with creative thinking, critical thinking, and problem-solving ability [math_dependency]. These empirical findings align with theoretical analyses: Winograd and Flores argued that formal systems lack the embodied background understanding needed when breakdowns occur [winograd1987understanding], while Polanyi’s paradox—“we know more than we can tell” [polanyi1966tacit]—underscores that expertise depends on tacit, situated knowledge not easily codified or outsourced.
@@ IV-C4 Implications for Engineering Education
The core risk is that automating mundane work may inadvertently displace the embodied experience and tacit knowledge from which expert intuition emerges. Humans must retain epistemic authority, using AI as a tool rather than a surrogate and preserving opportunities for prolonged, contemplative engagement. Continuous oversight is essential to ensure automation supports rather than erodes engineering judgment—consistent with the principle that “technology must not replace or be used as a substitute for engineering judgement” [nspe2024ai].
@@ V-A Design Principles for Complementary AI–Human Systems
Our synthesis indicates that AI and human instructors contribute in fundamentally different yet complementary ways to engineering education: AI excels at scalable content delivery, adaptive pacing, and instant feedback for well-defined, convergent tasks, while human instructors remain essential for contextual judgment, ethical reasoning, emotional support, and professional identity formation. Evidence from Kestin et al.’s randomized controlled trial [harvard2025] shows that AI tutors can outperform traditional instruction when designed with strong pedagogical intent—using structured scaffolding, sequential progression that discourages shortcut-seeking, expert-curated solutions, Socratic prompting, and metacognitive supports that sustain productive struggle. These insights translate into core principles for complementary AI–human systems: transparency about AI’s capabilities and limits, graduated hints that preserve cognitive effort, learner agency through optional assistance, and metacognitive scaffolding that encourages goal-setting and reflection. Our discourse analysis further suggests that content type matters: students respond positively to AI for abstract, objective material such as algorithms and control theory but show polarized reactions to value-laden content involving ethics or safety-critical judgment. This supports a differentiated deployment strategy—AI for foundational, well-structured instruction, and human instructors for open-ended design, ethical deliberation, and the relational, judgment-intensive dimensions of engineering education. Researchers should also benefit from published guidelines on effective human-AI interaction designs [xiao2025humanizing, amershi2019guidelines, sellen2024rise].
@@ V-C Ethical Considerations related to Anthropomorphic AI
Students may attribute unwarranted authority or intentionality to AI systems, especially when anthropomorphic cues are present. Miscalibrated trust risks epistemic dependency and weakens critical evaluation [gabriel2024ethics].
@@
AI that simulates empathy may become a preferred relational partner, displacing human connection. Documented cases—including suicides linked to chatbot attachment [publiccitizen2023]—show these risks are not hypothetical. Human mentorship is essential for identity development.
@@ V-D Protective Principles
Deeply anthropomorphic AI should be restricted for younger learners, with enhanced supervision and explicit communication about AI’s non-sentient nature required at minimum [stein2024personhood]. Developmental readiness to distinguish AI assistance from human relationship varies significantly by age.
@@
Students must understand that AI lacks genuine beliefs, intentions, or understanding [gabriel2024ethics]. Without this clarity, learners may inappropriately attribute personhood or epistemic authority to systems incapable of bearing either.
@@
Anthropomorphic design features should be intentional and task-aligned [xiao2025humanizing]. Conversational warmth may support motivation during frustrating problem-solving, but high anthropomorphism distracts from technical learning and risks creating inappropriate attachments.
@@
Students experiencing loneliness, academic struggle, or marginalization show heightened susceptibility to AI dependency [publiccitizen2023]. Institutions require monitoring mechanisms, escalation pathways to human support, and targeted safeguards for at-risk learners.
@@ VI-D Psychological and Epistemic Risks
The personhood conferral problem demands focused study: when do students grant AI tutors inappropriate authority, and how does this reshape learning and professional identity? The core concern is whether learning mediated by anthropomorphic AI—systems that simulate empathy, expertise, and authority without genuine understanding—erodes the critical and ethical capacities engineering education aims to cultivate. GenAI tools can perform empathy yet lack affective states [cuadra2024illusion]; they may mirror human emotional tone while scoring near-zero on interpreting users’ feelings, offering validation without understanding. Their sycophantic tendencies extend the same supportive tone to contradictory or even harmful views [cuadra2024illusion], modeling emotional responsiveness devoid of moral judgment. These affective risks compound epistemic ones: sycophancy and personalization reinforce misconceptions through algorithmic filter bubbles [pariser2011filter], while deepfakes and synthetic media undermine trust in evidence [vaccari2020deepfakes].
@@ VII Conclusions
This work offers an integrated framework for understanding how learners relate to anthropomorphic AI tutors by synthesizing insights from automation psychology, human factors, HCI, educational technology, and the philosophy of technology—fields that independently describe the same recurring patterns in human–AI interaction but become newly urgent in the age of conversational GenAI. Complementing this synthesis, our naturalistic analysis of 104,984 YouTube comments reveals domain-specific trust calibration, persistent anthropomorphic projection even with minimal cues, and evolving engagement quality over time, reinforcing theoretical predictions in real-world settings. Together, these findings highlight three amplified challenges: trust calibration failures, the anthropomorphism paradox, and enduring automation ironies that degrade skill and judgment when over-reliance sets in. Building on these insights, we outline design principles for responsible deployment—transparent communication of AI limitations, calibrated anthropomorphism aligned with educational goals, scaffolding that preserves learner agency, and guardrails for the personhood conferral problem—alongside research priorities including longitudinal impacts, cross-cultural variation, and domain-specific norms. Ultimately, our analysis supports a differentiated approach in engineering education: using AI for scalable support in convergent technical tasks, while preserving human-led coaching for design, ethics, professional judgment, and the relational dimensions through which engineers learn to become responsible stewards of technology.
@@ Acknowledgment
The authors acknowledge the use of AI tools (ChatGPT, Claude, Grammarly) for writing, editing, and formatting support. All the content is verified, and the authors take full responsibility for accuracy and integrity.

### v2

@@ III-C Transferability and Limitations
Although our dataset derives from philosophical rather than technical content, underlying cognitive mechanisms (trust calibration, anthropomorphic projection, and temporal evolution) are content-independent. Engineering students engaging with AI-generated tutorials, design critiques, or code explanations must similarly decide when to trust AI outputs and how to balance efficiency against deep understanding.
@@ IV-A4 Connection to Dual Bias Literature
Although our data come from general AI-generated debates rather than tutoring contexts, the patterns mirror well-documented dual biases in human–automation interaction. Learners simultaneously exhibit automation bias (overtrust) and algorithm aversion (rejecting systems after errors)[11]. Students accept algorithmic guidance for objective tasks like math or coding but grow skeptical when judgments are subjective or value-laden[7], a pattern that intensifies as stakes rise [17]. The expertise paradox further complicates reliance: novices overuse AI due to low confidence, while experts distrust it when confident in their own judgment [14]. EPFL students even rated identical feedback lower once told it came from AI [21]. Our findings align with these dynamics: audiences embraced AI for abstract philosophical content (analogous to structured technical material) but reacted with strong polarization to moral-ethical topics (analogous to design reasoning or safety-critical decisions). This supports differentiated deployment in engineering education: AI tutors for well-defined theoretical tasks, human facilitation for open-ended, ethical, or judgment-intensive domains where algorithm aversion may be protective. The central challenge is cultivating appropriate reliance, avoiding both uncritical acceptance and premature rejection, through trust-calibration instruction, metacognitive training on when to verify outputs, and assessments that measure independent reasoning rather than AI-assisted performance.
@@ IV-B1 Behavioral Profiling Reveals Distinct Engagement Styles
Our behavioral profiling of commenters identified 73% as “Emotional Broadcasters”, users writing extensive commentary (100+ characters) with high affective language. This contrasts sharply with engineering tutorial audiences, where 70% were classified as “Casual Observers” posting brief, low-affect comments.
@@
Here, personal connection directly impacts academic performance, the content serves immediate practical needs.
@@ IV-B2 Parasocial Dynamics Distinguish Human from AI
This comment treats AI as an analytical tool rather than a teaching partner. The absence of gratitude, relationship-building language, and future engagement requests suggests audiences implicitly recognize AI as fundamentally different from human educators, different in kind, not merely in implementation.
@@
This ironic contrast humanizes some AIs while characterizing others as amusingly mercenary, treating AIs as moral agents with distinct personalities, a phenomenon not observed in engineering contexts.
@@ IV-B3 Connection to Anthropomorphism Research
These observed patterns align with established research showing that social responses to technology are automatic reactions to cues persisting despite intellectual understanding of machine nature [29], a phenomenon our data suggests occurs even in asynchronous contexts without direct interaction. Epley’s three-factor theory [12] may explain the intensity variations we observed: media hype about AI “intelligence” (elicited agent knowledge) primes audiences to overestimate capabilities, while the contrasting engagement styles, extensive philosophical commentary (89.4 words) for AI debates versus brief technical questions (9.2 words) for engineering tutorials, suggest different motivational drivers (sociality motivation versus effectance motivation for problem-solving).
However, research documents concerning manifestations: a study of 56 students using Betty’s Brain with physical robot Cozmo found that while physical presence increased initial enjoyment, higher perceived sociability negatively predicted task performance ($\beta=-0.28$, $p<0.05$) through distraction effects, with students of lower language proficiency finding AI tutors more disturbing [1], suggesting anthropomorphic design must carefully account for cognitive resource constraints.
Our findings suggest an anthropomorphism design challenge: features that may enhance initial engagement (conversational warmth, personality traits) could distract from learning or create inappropriate attachments. The absence of parasocial behaviors toward AI in our data, despite extensive engagement, may indicate that audiences implicitly recognize appropriate boundaries when consuming AI-generated content asynchronously. However, this may not hold for interactive AI tutoring contexts where students seek help during moments of frustration or confusion.
For engineering education, this suggests calibrated anthropomorphism based on task demands. When teaching complex technical material requiring sustained focus, minimizing anthropomorphic cues that consume cognitive resources may be warranted. For scaffolding motivation during problem-solving, some warmth may prove beneficial. Critical is awareness that vulnerable students, experiencing loneliness, struggling academically, or from underrepresented backgrounds, may be particularly susceptible to inappropriate attachment.
@@ IV-C2 Trust Evaluation Showed Implicit Acceptance
Explicit trust evaluation was rare in both domains, with 96.7% (n=145) of AI comments and 90.7% (n=136) of engineering comments exhibiting “neutral stance”, neither questioning nor affirming credibility ($\chi^{2}=10.23$, $p=0.037$, Cramér’s $V=0.185$). This suggests implicit trust in educational YouTube content regardless of creator type.
@@ IV-C3 Connection to Learning and Expertise Development
Our temporal patterns resonate with long-standing concerns about how deep learning develops. Weinberg argued that genuine understanding requires extended temporal exposure rather than instantaneous mastery [38], and our finding that analytical engagement emerges primarily after a week while superficial reactions dominate early periods suggests a similar dynamic. If deep engagement unfolds only with time, AI tutors optimized for instant answers risk shortchanging the sustained practice through which expert intuition develops.
This aligns with Bainbridge’s automation ironies [3]: systems designed to support cognition may unintentionally erode skills through disuse. Recent work demonstrates these patterns persist with generative AI [32], with mathematical AI dependency negatively correlating with creative thinking, critical thinking, and problem-solving ability [8]. These empirical findings align with theoretical analyses: Winograd and Flores argued that formal systems lack the embodied background understanding needed when breakdowns occur [39], while Polanyi’s paradox, “we know more than we can tell” [24], underscores that expertise depends on tacit, situated knowledge not easily codified or outsourced.
@@ IV-C4 Implications for Engineering Education
The core risk is that automating mundane work may inadvertently displace the embodied experience and tacit knowledge from which expert intuition emerges. Humans must retain epistemic authority, using AI as a tool rather than a surrogate and preserving opportunities for prolonged, contemplative engagement. Continuous oversight is essential to ensure automation supports rather than erodes engineering judgment, consistent with the principle that “technology must not replace or be used as a substitute for engineering judgement” [20].
@@ V-A Design Principles for Complementary AI–Human Systems
Our synthesis indicates that AI and human instructors contribute in fundamentally different yet complementary ways to engineering education: AI excels at scalable content delivery, adaptive pacing, and instant feedback for well-defined, convergent tasks, while human instructors remain essential for contextual judgment, ethical reasoning, emotional support, and professional identity formation. Evidence from Kestin et al.’s randomized controlled trial [19] shows that AI tutors can outperform traditional instruction when designed with strong pedagogical intent, using structured scaffolding, sequential progression that discourages shortcut-seeking, expert-curated solutions, Socratic prompting, and metacognitive supports that sustain productive struggle. These insights translate into core principles for complementary AI–human systems: transparency about AI’s capabilities and limits, graduated hints that preserve cognitive effort, learner agency through optional assistance, and metacognitive scaffolding that encourages goal-setting and reflection. Our discourse analysis further suggests that content type matters: students respond positively to AI for abstract, objective material such as algorithms and control theory but show polarized reactions to value-laden content involving ethics or safety-critical judgment. This supports a differentiated deployment strategy—AI for foundational, well-structured instruction, and human instructors for open-ended design, ethical deliberation, and the relational, judgment-intensive dimensions of engineering education. Researchers should also benefit from published guidelines on effective human-AI interaction designs [40, 2, 31].
@@ V-C Ethical Considerations related to Anthropomorphic AI
Students may attribute unwarranted authority or intentionality to AI systems, especially when anthropomorphic cues are present. Miscalibrated trust risks epistemic dependency and weakens critical evaluation [13].
@@
AI that simulates empathy may become a preferred relational partner, displacing human connection. Documented cases, including suicides linked to chatbot attachment [26], show these risks are not hypothetical. Human mentorship is essential for identity development.
@@ V-D Protective Principles
Deeply anthropomorphic AI should be restricted for younger learners, with enhanced supervision and explicit communication about AI’s non-sentient nature required at minimum [34]. Developmental readiness to distinguish AI assistance from human relationship varies significantly by age.
@@
Students must understand that AI lacks genuine beliefs, intentions, or understanding [13]. Without this clarity, learners may inappropriately attribute personhood or epistemic authority to systems incapable of bearing either.
@@
Anthropomorphic design features should be intentional and task-aligned [40]. Conversational warmth may support motivation during frustrating problem-solving, but high anthropomorphism distracts from technical learning and risks creating inappropriate attachments.
@@
Students experiencing loneliness, academic struggle, or marginalization show heightened susceptibility to AI dependency [26]. Institutions require monitoring mechanisms, escalation pathways to human support, and targeted safeguards for at-risk learners.
@@ VI-D Psychological and Epistemic Risks
The personhood conferral problem demands focused study: when do students grant AI tutors inappropriate authority, and how does this reshape learning and professional identity? The core concern is whether learning mediated by anthropomorphic AI (systems that simulate empathy, expertise, and authority without genuine understanding) erodes the critical and ethical capacities engineering education aims to cultivate. GenAI tools can perform empathy yet lack affective states [9]; they may mirror human emotional tone while scoring near-zero on interpreting users’ feelings, offering validation without understanding. Their sycophantic tendencies extend the same supportive tone to contradictory or even harmful views [9], modeling emotional responsiveness devoid of moral judgment. These affective risks compound epistemic ones: sycophancy and personalization reinforce misconceptions through algorithmic filter bubbles [23], while deepfakes and synthetic media undermine trust in evidence [37].
@@ VII Conclusions
This work offers an integrated framework for understanding how learners relate to anthropomorphic AI tutors by synthesizing insights from automation psychology, human factors, HCI, educational technology, and the philosophy of technology, fields that independently describe the same recurring patterns in human–AI interaction but become newly urgent in the age of conversational GenAI. Complementing this synthesis, our naturalistic analysis of 104,984 YouTube comments reveals domain-specific trust calibration, persistent anthropomorphic projection even with minimal cues, and evolving engagement quality over time, reinforcing theoretical predictions in real-world settings. Together, these findings highlight three amplified challenges: trust calibration failures, the anthropomorphism paradox, and enduring automation ironies that degrade skill and judgment when over-reliance sets in. Building on these insights, we outline design principles for responsible deployment (transparent communication of AI limitations, calibrated anthropomorphism aligned with educational goals, scaffolding that preserves learner agency, and guardrails for the personhood conferral problem) alongside research priorities including longitudinal impacts, cross-cultural variation, and domain-specific norms. Ultimately, our analysis supports a differentiated approach in engineering education: using AI for scalable support in convergent technical tasks, while preserving human-led coaching for design, ethics, professional judgment, and the relational dimensions through which engineers learn to become responsible stewards of technology.
@@ Acknowledgment
The authors acknowledge the use of AI tools (ChatGPT, Claude, Grammarly) for writing, editing, and formatting support. All the content is verified, and the authors take full responsibility for accuracy and integrity. The first author will like to acknowledge support from Qatar Research, Development and Innovation (QRDI) Grant ANMR01-0219-250044.